OpenStudy (kitkat16):

Please help writing the expression of 1+5+25+125+625

OpenStudy (some.random.cool.kid):

what do you mean

OpenStudy (anonymous):

the answer is 781. you add them

OpenStudy (kitkat16):

\[\sum_{i}^{5?}=0^5^i or \sum_{i}^{5?}=1^5^i\]

OpenStudy (anonymous):

Each one is multiplied by 5... what do you mean by writing the expression

OpenStudy (some.random.cool.kid):

so if your trying to add then he/she is correct

OpenStudy (kitkat16):

not sure how to use the equation thing I don't think it shows up

OpenStudy (some.random.cool.kid):

but what do you mean

OpenStudy (kitkat16):

Ill post a pic

OpenStudy (some.random.cool.kid):

ok

OpenStudy (kitkat16):

I think it is the last one

OpenStudy (some.random.cool.kid):

yeah maybe

OpenStudy (some.random.cool.kid):

im still reading it

OpenStudy (anonymous):

lol random

OpenStudy (some.random.cool.kid):

looks right

OpenStudy (anonymous):

speaking of which book are you reading random? >.<

OpenStudy (kitkat16):

No the one i choose was wrong can anyone explain this to me? @shinebrightlikeadimon

OpenStudy (kitkat16):

@hartnn if I was to write this equation is it is the first 5 terms is it

OpenStudy (kitkat16):

I believe it to be c.

hartnn (hartnn):

1+5+25+125+625 = 5^0 +5^1+5^2 +5^3 +5^4 \(\Large \sum \limits_{i=0}^4 5^i \)

hartnn (hartnn):

A

OpenStudy (kitkat16):

Thank you I figured it out yesterday.

hartnn (hartnn):

welcome ^_^